OpenStudy (anonymous):

How many numbers between 0 and 300 are divisible by 3 but not by 5 or 7? If possible send the explanation also Thanks in advance

OpenStudy (anonymous):

Well there are 100 numbers divisible by 3 in the set 0 to 300. Of those, 300/(3*5)=20 are divisible by 3 and 5. Of the remaining, 300/(3*7)=14 are divisible by 3 and 7. However, 35*3n is already counted which means 105,210 (2 numbers). So I think the answer would be 100-20-14+2=68. I could absolutely be wrong but it seems logical at the time. ;p

OpenStudy (anonymous):

I would agree with @JonathanHocker

OpenStudy (anonymous):

\[ \left\lfloor \frac {300}{3}\right\rfloor - \left(\left\lfloor \frac {300}{3\times5}\right\rfloor + \left\lfloor \frac {300}{3\times7}\right\rfloor - \left\lfloor \frac {300}{3\times 5\times 7}\right\rfloor\right) \]

OpenStudy (anonymous):

What are those brackets? Some sort of floor function?